You have purchased merchandise with a list price of USD 36,000. Because you are a wholesaler,

you are granted a trade discount of 49.6 per cent. The cash discount terms are 2/EOM, n/60. How much will you

remit if you pay the invoice by the end of the month of purchase? How much will discounts on payment you remit if

you do not pay the invoice until the following month?


1
Expert's answer
2021-07-27T02:00:02-0400

The amount to remit will be the list price less trade discount to get the invoice price, which you less the cash discount to get the final amount paid. This is calculated as follows:

List price = 36,000

Trade discount = 49.6"\%"

Invoice price = List price "\times" (1 – Trade discount "\%" ) = 36,000 "\times" (1 – 49.6"\%" ) = USD 18,144

 

2/EOM means that a cash discount of 2% is allowed if you pay by the end of the month, while n/60 means the entire invoice is due 60 days from the invoice date.

Therefore, both the trade discount and the cash discount will apply if payment is made by the end of the month of purchase.

 

Amount to remit = Invoice price "\times" (1 – Cash discount "\%" ) = 18,144 "\times" (1 – 2"\%" ) = USD 17,781.12

The amount to be remitted if you pay the invoice by the end of the month of purchase = USD 17,781.12

 

The discount on the payment you remit if you do not pay the invoice until the following month will only be the trade discount. The cash discount will not apply in this case since the amount will be paid beyond the end of the month for which the cash discount applies.


Amount of discount = List price "\times" Trade discount "\%" = 36,000 "\times" 49.6"\%" = $17,856

Amount of discount = USD 17,856